OpenStudy (anonymous):

lim_{rightarrow infty}\frac{ sqrt{4n^3+5} }{ 6n^3+7 } i need this one for tommorow pls

OpenStudy (anonymous):

We can't see it...

OpenStudy (anonymous):

\[\lim_{\rightarrow \infty}\frac{ \sqrt{4n^3+5} }{ 6n^3+7 } \]

OpenStudy (anonymous):

i need to find the limit

OpenStudy (anonymous):

\[\left[ \frac{ \infty }{ 0 } \right]\]

OpenStudy (anonymous):

should end like this

OpenStudy (anonymous):

What you have is a case of infinity/infinity. This fits into l'Hotial's condition.

OpenStudy (anonymous):

nah, just look denominator is degree 3, numerator is degree \(\frac{3}{2}\) denominator is larger, limit is zero